OpenStudy (anonymous):

Group the terms with variables on one side of the equal sign, and simplify. 39y = 4y + 11

OpenStudy (anonymous):

\(39y = 4y +11\) \(39y - 4y = 11\) \(35y =11\) \(y = \frac{11}{35}\)

OpenStudy (anonymous):

Group the terms with variables on one side of the equal sign, and simplify. 5z + 3 = 36z 3 = 41z 8z = 36 3 = 31z −3 = 31z

OpenStudy (anonymous):

Close! But you added \(5z\) to both sides at the first step instead of subtracting it. So it's: \(5z+3 = 36z\) \(3 = 36z - 5z\) \(3 = 31z\) Can you take it from there?

OpenStudy (anonymous):

What is the value of \(z\) then?

OpenStudy (anonymous):

can you show me step by step how to

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The first step to solve for the variable (in our case, \(z\)) is to get the similar \(z\) terms on one side of the equation: So take \(5z + 3 = 36z\) and we want to have all the terms with "\(z\)" on one side, so lets subtract "\(5z\)" from both sides. \(5z + 3 - 5z = 36z - 5z\) which reduces to \(3 = 31z\). Do you understand? Also if we wanted to solve for \(z\) we want to divide both sides by \(31\). This is because we don't want the value of "\(31z\)", we want the value of "\(z\)"! So \(z = \frac{3}{31}\).